Solve the system by the elimination method.
0.3x-0.2y=4
0.4x+0.5y=-55/17
Multiply equations by 10 to eliminate decimals.
3x-2y=40
4x+5y= -550/17
Then to eliminate y multiply the first equation by5 and the second equation by 2.
15x-10y=200
8x+10y=-1100/17
I have no idea what do do from here...Everytime I try my fractions come out wrong and I get lost.

Answer by checkley77(12844)

0.3x-0.2y=4
0.4x+0.5y=-55/17
----------
3x-2y=40
4x+5y= -550/17
------------
15x-10y=200
8x+10y=-1100/17 now add these 2 equations.
------------
23x=200-1100/17
23x=(200*17-1100)/17
23x=(3400-1100)/17
23x=2300/17
x=(2300/17)(1/23)
x=100/17 answer.
.3*100/17-.2y=4
30/17-.2y=4
-.2y=4-30/17
-.2y=(4*17-30)/17
-.2y=(68-30)/17
-.2y=38/17
y=(38/17)(1/-.2)
y=38/-3.4
y=-19/1.7
y=-190/17